OpenStudy (anonymous):

find y'(1) if y = tan^-1 x + Sec^-1 sqrt(1+x^2) a. 0 b.2 c.1 d.3

OpenStudy (anonymous):

1

OpenStudy (anonymous):

PUT TAN X =Z

OpenStudy (anonymous):

YOU WILL GET Y = 2 * TAN INVERSE OF X

OpenStudy (anonymous):

what is z?

OpenStudy (anonymous):

Z IS ANY ANGLE. IF U PUT X=Z U WILL GET Y= Z+Z BUT Z= TAN^-1 (X) HENCE Y=2* TAN^-1(X) Y'=2 * 1/(1+X^2) AND Y(1)=1